OpenStudy (anonymous):

Am I right? Choose the correct description of the graph of the inequality -x greater than or equal to 3. Closed circle on 3, shading to the right.

OpenStudy (anonymous):

\[-x\geq3\] \[x\leq-3\]

OpenStudy (anonymous):

was that a minus sign or a dash?

OpenStudy (anonymous):

\[-x \ge 3\]

OpenStudy (anonymous):

ok then change the sign of both sides and change the sense of the inequality, you get \[x\leq-3\]

OpenStudy (anonymous):

don't forget you are solving for x, not "-x"

OpenStudy (anonymous):

So, then it would be: Closed circle on -3, shading to the left;?

OpenStudy (anonymous):

therefore it is the interval to the left of -3 yes, you are right, closed at -3, shaded left

OpenStudy (anonymous):

Thank you! :DD

OpenStudy (anonymous):

yw!